How much energy (heat) is required to convert 35.8 g of H2O(l) at 42.0°C to H2O(g) at 105.0°C?

Specific heat water = 4.18 J/g/oC

Specific heat steam = 1.84 J/g/oC

ΔHvap (water) = 40.7 kJ/mol
1
Expert's answer
2020-10-26T14:54:05-0400

Solution:

Three calculations, then add the results of those calculations for the final answer.

1) Energy required to warm the water from 42.0 C to 100.0 C

2) Energy required to boil the water at 100.0 C

3) Energy required to warm the steam from 100.0 C to 105.0 C

-------------------

"Q_{1}=(35.8g)(58^{o} C)(4.184( \\tfrac{J}{g*C}\t))=8687.65J = 8.68kJ"


"Q_{2}=( \\dfrac{35.8g}{18( \\tfrac{g}{mol}\t)}\t)(40.7( \\tfrac{kJ}{mol}\t)=80.94 kJ"


"Q_{3}=(35.8g)(5^{o}C)(1.84 ( \\tfrac{J}{g*^{o}C}\t))=329.36J=0.329kJ"


"Q_{total}=Q_1+Q_2+Q_3=8.68+80.94+0.329=90kJ"


Answer:


"Q_{total} = 90kJ"